OpenStudy (anonymous):

if f(x)=[[x-3.2]]-1/2, what are the values of f(5.8) and (8.5), respectively. 1 and 2 0.5 and 3.5 1.8 and 2.15 1 and 2.5

OpenStudy (dtan5457):

Your answers are wrong.

OpenStudy (dtan5457):

f(x)=[[x-3.2]]-1 /2 This is f(x)-4.2, then divided by 2. f(8.5) is in one of your answer choices 8.5-4.2=4.3/2=2.15 however (5.8) is not..

OpenStudy (dtan5457):

check for typos..

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

do you think you can help me with graphing linear equations?